# Hybrid User Interfaces: Past, Present, and Future of Complementary Cross-Device Interaction in Mixed Reality

### v1

@@ 5.2 Use Cases
We analyzed usage scenarios as described by the papers in our corpus and further clustered them into common use cases. The most prominent use case was the area of visual analytics 26, ranging from abstract visualization for immersive analytics [25] (e.g., [67, 27, 119, 124, 18, 87]), dashboards[144], user study analysis  [66, 103], and scientific visualizations [13, 2, 96]. This is unsurprising, since concepts such as multiple-coordinated views are widely established and lend themselves well to configurations such as the asymmetric mirror. This also partially overlaps with the medical 4 domain, which uses HUIs for 3D examination [13, 2, 96] and surgery  [72]. Another popular use case is 3D modeling 11 [139, 118, 117], in which we include 3D sketching [3, 34]. Several records present the area of development and authoring 7, including development toolkits [51, 103, 114] or programming-related tasks [120, 8, 145, 123]. We attribute several records to general productivity 11 tools, such as window management [47], extending desktop configurations [110, 109, 80], file transfer [133], or general user interface improvements [81, 15, 78]. We also found several records in domains such as gaming 4 [143, 101], entertainment 4 (e.g., music [40, 83], television [79, 5]), collaboration 6 [19, 121, 49, 130], or text entry 5 [55, 85, 7]. Lastly, the remainder had study-specific 6 or other 11 use cases.
@@ 5.3 Devices and Combinations
As publications in our corpus combine multiple interaction components (i.e., devices) in a complementary way, we recorded device technologies and their most common combinations (see Fig. 6). For records that did not have specific terminology to describe their hardware (e.g., [141]), we used current terms based on the dimensions of the device.
In terms of MR interaction components, AR HWDs 63 were the most common device technologies, with a large part of systems using OST AR HWDs 50 and others using VST AR HWDs 13. For handheld AR 11, we differentiate between handheld AR on smartphones 6 and handheld AR on tablets 5. Other types of AR systems were less common, such as projector-based spatial AR 4, stereoscopic projections 2, and CAVEs 2. Lastly, since we also included VR environments, our corpus contained systems with VR HWDs 13.
For 2D interaction components, we observed a similar spread between available device technologies: Several systems used mobile platforms 55 such as smartwatches 5, smartphones 21, tablets 22, and laptops 7. Stationary interaction components 48 were almost as common, including desktops 24, projectors 6, large displays 8 (e.g., wall-displays), and tabletops 10.
Looking at the device combinations (see Fig. 6 (left)) over the years, we can observe an increase in device variation as more new form factors become available. This proliferation has resulted in a variety of device combinations (see Fig. 6 (right)), with a focus on HWDs, although other MR devices were also used. Most often, 2D interaction components were represented by handheld devices (e.g., smartphones), or desktop interaction components. The usage of AR OST HWDs was preferred for the described 2D interaction components. For smartwatches, there is a preference for AR VST HWDs 4, probably because their field of view is more suitable for this form factor (cf. [104]).
@@ 5.4 Contribution Types
We classified papers in our corpus following the definition of Wobbrock and Kientz [152] (see Fig. 7). As some papers provide more than one contribution (e.g., an artifact that was used as an apparatus in an empirical user study), we distinguish between primary and secondary contributions and report the results accordingly (i.e., 15 / 4).
Artifact 56 / 8
@@
Two papers in our corpus [87, 158] are each categorized with a single primary contribution and two equivalent secondary contributions: Both present a design space as their main contribution (theory), followed by a system (artifact) used as the apparatus of a user study (empirical).
No papers were classified as datasets, surveys, or opinion contributions (see Fig. 7). This is not surprising, as we focused on actual systems, which resulted in most artifact and empirical contributions. Similarly, some of the papers that could fulfill the criteria of other types of contribution were not considered, as they typically did not meet the eligibility criteria of our survey. Five papers contributed a theoretical contribution (i.e., a design space [87, 158]), showing that design spaces can be a way of exploring this nascent research area without necessarily implementing an extensive system or study apparatus.
@@ 5.5 Evaluation Strategies
For papers that provided a primary or secondary empirical contribution, we coded their evaluation strategies as proposed by Ledo et al. [88] (see Fig. 7). Most papers featured one empirical contribution, others combined multiple user studies with different strategies.
The most common evaluation strategies within our corpus were empirical user studies that examine how users interact with a system 55. In their cross-device taxonomy, Brudy et al. [16] further split this evaluation strategy into qualitative and quantitative usage and informative (observational and elicitation) evaluations: While usage 37 focuses on the usability and usefulness of the system and how it is appropriated [16], informative 18 evaluations involve studies that precede and inform the development of a system, involving users in the design process [16]. Demonstrations 16 are used to describe how systems are employed in an actual use case scenario but do not necessarily involve a real system implementation. In contrast, technical performance 4 evaluations (cf. technical evaluation [16]) focus on benchmarking an implemented system in terms of its technical capabilities. Lastly, several systems did not include any kind of evaluation 18. Similarly, we did not observe any kind of heuristics 0 evaluation (i.e., using guidelines to analyze usability), which could be attributed to the lack of appropriate guidelines for MR-HUIs.
Figure 7: Distribution of primary contribution types and evaluation strategies over time within our corpus (years without records were omitted). Here, each cell within a row (i.e., year) is shaded based on how many papers presented this contribution or evaluation type in the respective year, visualizing the relative development over time (i.e., in a column).
@@ 5.6 Summary of Insights
Term Fragmentation.  Our corpus shows that terminology in prior research is mostly evenly split between hybrid user interfaces 22 and cross-device interaction 17, while the emergent area of cross-reality 8 is gaining traction. A large number of eligible records 35 in this survey were found in adjacent research areas, barely mentioning such hybrid device combinations and thus making it difficult to search for and identify relevant prior work. This fragmentation is also reflected in their specificity, ranging from rather broad (e.g., cross-device interaction 17, multi-device interaction 10) to very narrow (e.g., hybrid ⟨other⟩ 16) descriptions. While broad terms encapsulate a breadth of unrelated systems, overly specific denominators may lead to fragmentation and could impede the understanding of the field.
@@
A similar fragmentation can be seen in the use of terminology to describe design dimensions. Although the cross-device taxonomy [16] provides an appropriate framework, we adapted many dimensions to better capture the unique design dimensions of MR-HUIs. Many useful terms that apply to MR-HUIs are hidden within artifact contributions (e.g., anchoring [114], dependency [158], configurations such as augmented display [118, 120], and VESADs [104]).
@@
Parallel usage is predominant.  We found that parallel usage of multiple interaction components is by far the most predominant design choice. We think the reasons are twofold. (1) Parallel usage offers greater design possibilities. In contrast, exclusive (and to some extent serial) usage considers one interaction component at a time, thus limiting the design potential. (2) As the time between using different interaction components exclusively increases, it becomes harder to see the whole system as one coherent interface—and harder to classify. Systems with exclusive usage might be better described as cross-reality, which is concerned with the general usage of multiple systems in different “realities”.
@@
Optical see-through is the prevalent hardware choice, despite its drawbacks.  AR HWDs 63 were the most common MR interaction component. Although current VST HWDs 13 offer a wider field of view for digital content [104], the use of OST HWDs 50 was much more prevalent. We attribute this to (1) the unrestricted real-world field of view of OST HWDs greatly facilitating interaction with 2D interaction components; (2) using OST HWDs further emphasizes the complementary nature of HUIs, as the addition of a secondary interaction component offsets the drawbacks of the AR HWD; and (3) VST HWDs having only recently matured enough to be used in conjunction with other devices (e.g., in terms of text legibility due to limited pass-through resolution). We also observed a steady increase in empirical contributions over the past years, indicating that the hardware is now mature enough to conduct studies that are not confounded by hardware restrictions.
@@
Lack of collaborative systems.  Our corpus shows a distinct lack of multi-user systems 19, especially within public spaces 0. Although there is a great deal of work in the field of computer-supported cooperative work (CSCW) for collaboration across asymmetric devices—which we intentionally excluded from our survey—we see great potential in the use of MR-HUIs in collaborative scenarios.
@@ 6 Key Challenges and Research Agenda
6.1 Transitioning Between Interaction Components
@@
HUIs draw their strength from the combination of complementary interaction components. While this can be beneficial, it also introduces several challenges concerning visual attention switching, transferring content between interaction components, and perceptual content synchronization.
@@
Visual attention switching.  Introducing multiple potential visual output components divides the user’s attention among each component, increasing mental load [111]. Prior work has investigated such visual attention switches in multi-display environments [112] and proposed several metrics to potentially reduce strain. Although MR-HUIs can benefit from these insights, their design space is much broader in terms of display layout, enabling display layouts that are unbound by reality. Still, current hardware limitations have to be considered, as several works [35, 54, 104, 151] opt for VST AR HWDs to avoid differences in focal planes, which can further increase mental load [35]. Overall, more research is required to investigate potential effects and solutions for different display layouts [82].
@@
Transferring content between interaction components.  MR-HUIs mix 2D and 3D content to make use of each interaction component’s strengths. While the distribution is often static (e.g., determined by the designer of the system), several systems have demonstrated the utility of transferring content from a 2D interaction component to the MR environment [131, 153, 148, 2]. Specifically, prior work by Schwajda et al. [131] highlights the importance of animations and linking when “pulling” content from a 2D interaction component into an MR environment (cf. showing Bézier curves to selected MR content [67, 144]). Furthermore, recent research in the area of immersive analytics investigated visualization transformations between 2D and 3D [89, 90]. Several transition metaphors have been explored in prior work (e.g., “pulling” sticky notes from a smartphone [153]), but further investigations are necessary to standardize and evaluate these techniques [131].
@@
Perceptual synchronization.  Another challenge is found within the mapping and synchronization of content and interaction across components. While prior research has shown that an asymmetry of interaction (e.g., decoupling spatial interaction [132]) or information (e.g., showing simplified 2D views of 3D content [71, 73]) can be beneficial, this may come at an increased cost of user perception: For example, how does a 2D interaction affect its 3D equivalent, and how can we communicate this asymmetry?
@@
Aside from these conceptual challenges in establishing a consistent mental model, we also have to consider technological challenges. In this context, prior work found that cross-device interaction techniques can be highly sensitive to network latency [92]. This sensitivity may be even more pronounced in MR-HUIs, especially if they appear as one conceptual device (e.g., the virtual screen in a VESAD configuration may lag behind the real screen). Here, more research is necessary to investigate how inevitable technological factors (e.g., latency) can confound findings.
6.2 Unchaining Interaction Component Capabilities
@@
The past decade has seen a surge in both device variety (e.g., smartwatches, smartrings) and device capabilities (e.g., inside-out spatial tracking), which can also be seen in the increase in interaction component combinations. Yet, 2D interaction components such as smartphones are restricted by their form factor, favoring device ergonomics and portability but limiting their potential output capabilities. By combining them with MR interaction components (e.g., AR HWDs), we can “unchain” their capabilities, enabling entirely new interaction possibilities (e.g., offloading menu items [117, 104, 15]), thereby inching closer to a form of “universal interaction” [82] and ultimately, Weiser’s vision of ubiquitous computing [149]. While this can further expand the design space, we also recognize the need to establish guidelines to better understand the trade-offs and potential of each device (cf. [146]).

### v2

@@ 5.2 Use Cases
We analyzed usage scenarios as described by the papers in our corpus and further clustered them into common use cases. The most prominent use case was visual analytics 33, ranging from abstract visualization for immersive analytics [chandler2015immersive] (e.g., [hubenschmid2021stream, chulpongsatorn2023holotouch, reipschlager2021personal, satkowski2021insitu, butscher2018clusters, langner2021marvis]), dashboards[vock2021idiar], user study analysis  [hubenschmid2022relive, nebeling2020mrat], and scientific visualizations [bornik2006hybrid, aigner2023cardiac, luo2021exploring]. This is unsurprising, since concepts such as multiple-coordinated views are widely established and lend themselves well to configurations such as the asymmetric mirror. This partially overlaps with the medical 7 domain, which uses HUIs for 3D examination [bornik2006hybrid, aigner2023cardiac, luo2021exploring] and surgery  [ilie2004combining]. Another popular use case is 3D modeling 13 [surale2019tabletinvr, reipschlager2020augmented, reipschlager2019designar], including 3D sketching [arora2018symbiosissketch, drey2020vrsketchin]. Several records present the area of development and authoring 13, including development toolkits [gonzalez2024xdtk, nebeling2020mrat, reichherzer2021secondsight] or programming-related tasks [reipschlager2018debugar, berns2019myr, wang2014coordinated, lunding2025arthur]. We attribute several records to general productivity 16 tools, such as window management [feiner1991hybrid], extending desktop configurations [pavanatto2021we, pavanatto2024multiple, kim2023perspective], file transfer [serrano2015gluey], note-taking [qiu2025marginalia], or general user interface improvements [knierim2021smartphone, brasier2021arenhanced, kane2009bonfire]. We also found several records in domains such as gaming 5 [tolstoi2015towering, mueller2015capture], entertainment 6 (e.g., music [elvezio2018hybrid, kohen2020mixr], television [kawakita2014augmented, baillard2017multidevice]), collaboration 6 [butz1999enveloping, rekimoto1999augmented, friedl-knirsch2023exploring, schroder2023collaborating], or text entry 6 [grubert2023text, krug2022clear, bang2023enhancing]. Lastly, the remainder had study-specific 21 or other 19 use cases.
@@ 5.3 Devices and Combinations
As publications in our corpus combine multiple IO components (i.e., devices) in a complementary way, we recorded device technologies and their most common combinations (see Fig. 6). For records that did not have specific terminology to describe their hardware (e.g., [szalavari1997personal]), we used current terms based on the dimensions of the device.
In terms of MR IO components, AR HWDs 87 were the most common device technologies, with a large part of systems using OST AR HWDs 69 and others using VST AR HWDs 22. For handheld AR 14, we differentiate between handheld AR on smartphones 8 and handheld AR on tablets 8. Other types of AR systems were less common, such as projector-based spatial AR 5, stereoscopic projections 2, and CAVEs 3. Lastly, since we also included VR environments, our corpus contained systems with VR HWDs 22.
For 2D IO components, we observed a similar spread between available device technologies: Several systems used mobile platforms 66 such as smartwatches 8, smartphones 33, tablets 29, and laptops 9. Stationary IO components 59 were almost as common, including desktops 37, projectors 6, large displays 9 (e.g., wall-displays), and tabletops 10.
Looking at the device combinations (see Fig. 6 left) over the years reveals an increase in device variation as more new form factors become available. This proliferation resulted in a variety of device combinations (see Fig. 6 right), with a focus on HWDs, although other MR devices were also used. Most often, 2D IO components were represented by handheld devices (e.g., smartphones), or desktop IO components. The usage of AR OST HWDs was preferred for the described 2D IO components.
@@ 5.4 Contribution Types
We classified papers in our corpus following the definition of Wobbrock and Kientz [wobbrock2016research]. As some papers provide more than one contribution (e.g., an artifact that was used as an apparatus in an empirical user study), we distinguish between primary and secondary contributions and report the results accordingly (i.e., 15 / 4).
Artifact 73 / 11
@@
Dataset 1 / 0
contributions offer a curated corpus for a specific topic to help design future MR-HUIs and layouts.
Two papers in our corpus [langner2021marvis, zhu2020bishare] are each categorized with a single primary contribution and two equivalent secondary contributions: Both present a design space as their main contribution (theory), followed by a system (artifact) used as the apparatus of a user study (empirical).
No papers were classified as surveys or opinion contributions. This is not surprising, as we focused on actual systems, which resulted in most artifact and empirical contributions. Similarly, some of the papers that could fulfill the criteria of other types of contribution were not considered, as they typically did not meet the eligibility criteria of our survey. Five papers contributed a theoretical contribution (i.e., a design space [langner2021marvis, zhu2020bishare]), showing that design spaces can be a way of exploring this nascent research area without necessarily implementing a system or study apparatus.
@@ 5.5 Evaluation Strategies
For papers that provided a primary or secondary empirical contribution, we coded their evaluation strategies [ledo2018evaluation]. Most papers featured one empirical contribution; others combined multiple user studies with different strategies.
The most common evaluation strategies within our corpus were empirical user studies that examine how users interact with a system 73. In their cross-device taxonomy, Brudy et al. [brudy2019crossdevice] further split this evaluation strategy into qualitative and quantitative usage and informative (observational and elicitation) evaluations: While usage 59 focuses on the usability and usefulness of the system and how it is appropriated [brudy2019crossdevice], informative 23 evaluations involve studies that precede and inform the development of a system, involving users in the design process [brudy2019crossdevice]. Demonstrations 21 are used to describe how systems are employed in an actual use case scenario but do not necessarily involve a real system implementation. In contrast, technical performance 5 evaluations (cf. technical evaluation [brudy2019crossdevice]) focus on benchmarking an implemented system in terms of its technical capabilities. Lastly, several systems did not include any kind of evaluation 21. Similarly, we did not observe any kind of heuristics 0 evaluation (i.e., using guidelines to analyze usability), which could be attributed to the lack of appropriate guidelines for MR-HUIs.
@@ 5.6 Summary of Insights
Term Fragmentation.  Our corpus shows that terminology in prior research mostly favors the broader term cross-device interaction 65 over hybrid user interfaces 29, while the emergent area of cross-reality 14 is gaining traction. A large number of eligible records in this survey were found in other research areas, barely mentioning such hybrid device combinations and thus making it difficult to search for and identify relevant prior work. This fragmentation is also reflected in their specificity, ranging from rather broad (e.g., cross-device interaction 26, multi-device interaction 10) to very narrow (e.g., hybrid ⟨other⟩ 20) descriptions. While broad terms encapsulate a breadth of unrelated systems, overly specific denominators may lead to fragmentation and could impede the understanding of the field.
@@
A similar fragmentation can be seen in the use of terminology to describe design dimensions. Although the cross-device taxonomy [brudy2019crossdevice] provides an appropriate framework, we adapted many dimensions to better capture the unique design dimensions of MR-HUIs. Many useful terms that apply to MR-HUIs are hidden within artifact contributions (e.g., anchoring [reichherzer2021secondsight], dependency [zhu2020bishare], configurations such as augmented display [reipschlager2020augmented, reipschlager2018debugar], and VESADs [normand2018enlarging]).
@@
Parallel usage is predominant.  We found that parallel usage of multiple IO components is by far the most predominant design choice. We think the reasons are twofold. (1) Parallel usage offers greater design possibilities. In contrast, exclusive (and to some extent serial) usage considers one IO component at a time, thus limiting the design potential. (2) As the time between using different IO components exclusively increases, it becomes harder to see the whole system as one coherent interface—and harder to classify. Systems with exclusive usage might be better described as cross-reality, which is concerned with the general usage of multiple systems in different “realities”.
@@
Optical see-through is the prevalent hardware choice, despite its drawbacks.  AR HWDs 87 were the most common MR IO component. Although current VST HWDs 22 offer a wider field of view for digital content [normand2018enlarging], the use of OST HWDs 69 was much more prevalent. We attribute this to (1) the unrestricted real-world field of view of OST HWDs greatly facilitating interaction with 2D IO components; (2) using OST HWDs further emphasizes the complementary nature of HUIs, as the addition of a secondary IO component offsets the drawbacks of the AR HWD; and (3) VST HWDs having only recently matured enough to be used in conjunction with other devices (e.g., in terms of text legibility due to limited pass-through resolution). We also observed a steady increase in empirical contributions over the past years, indicating that the hardware is now mature enough to conduct studies that are not confounded by hardware restrictions.
@@
Lack of collaborative systems.  Our corpus shows a distinct lack of multi-user systems 21, especially within public spaces 2. Although there is a great deal of work in the field of computer-supported cooperative work (CSCW) for collaboration across asymmetric devices—which we intentionally excluded from our survey—we see great potential in the use of MR-HUIs in collaborative scenarios.
5.7 Potential and Pitfalls of MR-HUIs
MR-HUIs are an intentional combination of off-the-shelf devices that would otherwise be used on their own. This is one of their strengths: Instead of requiring bespoke hardware tailored to a narrow range of tasks (e.g., tangibles), MR-HUIs leverage the user’s existing device ecology.
However, this can also have several pitfalls: Users’ preferences for modalities might not necessarily reflect the design of the MR-HUI, as they might, for example, favor efficiency over effectiveness (cf. [zagermann2017memory]). In addition, as off-the-shelf devices are designed for general-purpose use, they inevitably involve trade-offs. For example, a smartphone’s screen may remain mostly unused inside MR environments, as users focus their visual attention on the content of the MR HWD. Should we favor specialized technologies, such as bespoke input devices, over general-purpose ones?
While specialized technologies may offer better performance for their intended use case, they lack the versatility to be used outside of their designated environments, leaving users with a plethora of single-purpose devices. In contrast, MR-HUIs leverage devices that are already within the user’s ecology, connecting not only hardware capabilities but also software ecosystems. This could increase the adoption of immersive technologies, as they can be integrated to enhance current workflows and practices, rather than supersede them. Even outside of the context of MR-HUIs, these devices remain valuable, as smartphones and MR HWDs, for example, can still be used on their own.
MR-HUIs thus represent a pragmatic design solution that leverages the complementarity of currently available device technologies within existing device ecologies. Even as hardware limitations become less relevant due to increasingly sophisticated MR HWDs, MR-HUIs can inform the design of future interaction techniques and configurations. For example, recent advances in MR HWDs may render migratory configurations obsolete, as users can simply simulate a desktop in MR. Yet, using MR-HUIs, we can already investigate potential techniques and devices (e.g., hybrid input devices [hubenschmid2025spatialmouse, hubenschmid2025revisiting]) that transcend the use of MR-HUI.
@@ 6 Key Challenges and Research Agenda
6.1 Transitioning Between IO components
@@
HUIs draw their strength from the combination of complementary IO components. While this can be beneficial, it also introduces several challenges concerning visual attention switching, transferring content between IO components, and perceptual content synchronization.
@@
Visual attention switching.  Introducing multiple visual output components divides the user’s attention among each component, increasing mental load [rashid2012cost]. Prior work has investigated such visual attention switches in multi-display environments [rashid2012factors] and proposed metrics to potentially reduce strain. Although MR-HUIs can benefit from these insights, their design space is much broader in terms of display layout, enabling display layouts that are unbound by reality. Still, factors such as distance [bang2025areading] or current hardware limitations have to be considered, as several works [eiberger2019effects, grubert2015multifi, normand2018enlarging, wieland2024push2ar] opt for VST AR HWDs to avoid differences in focal planes, which can further increase mental load [eiberger2019effects]. Overall, more research is required to investigate potential effects and solutions for different display layouts [knierim2023universal].
@@
Transferring content between IO components.  MR-HUIs mix 2D and 3D content to make use of each IO component’s strengths. While the distribution is often static (e.g., determined by the designer of the system), several systems have demonstrated the utility of transferring content from a 2D IO component to the MR environment [schwajda2023transforming, wu2020megereality, wang2024user, aigner2023cardiac]. Specifically, prior work [schwajda2023transforming] highlights the importance of animations and linking when “pulling” content from a 2D IO component into an MR environment (cf. showing Bézier curves to selected MR content [hubenschmid2021stream, vock2021idiar]). Furthermore, recent research in the area of immersive analytics investigated visualization transformations between 2D and 3D [lee2022design, lee2023deimos]. Several transition metaphors have been explored [wu2020megereality, zhao2025spatialtouch] and compared [rau2025traversing, cools2025comparison] in prior work, but further investigations are necessary to standardize these techniques.
@@
Perceptual synchronization.  Another challenge is found within the mapping and synchronization of content and interaction across components. While prior research has shown that an asymmetry of interaction (e.g., decoupling spatial interaction [sereno2022hybrid]) or information (e.g., showing simplified 2D views of 3D content [hollerer1999exploring, iquiapaza2024dear]) can be beneficial, this may come at an increased cost of user perception: For example, how does a 2D interaction affect its 3D equivalent, and how can we communicate this asymmetry?
@@
Aside from these conceptual challenges in establishing a consistent mental model, we also have to consider technological challenges. In this context, prior work found that cross-device interaction techniques can be highly sensitive to network latency [liu2023defining]. This sensitivity may be even more pronounced in MR-HUIs, especially if they appear as one conceptual device (e.g., the virtual screen in a VESAD configuration may lag behind the real screen). Here, more research is necessary to investigate how inevitable technological factors (e.g., latency) can confound findings.
6.2 Unchaining Device Capabilities
@@
The past decade has seen a surge in both device variety (e.g., smartwatches, smartrings) and device capabilities (e.g., inside-out spatial tracking), which can also be seen in the increase in IO component combinations. Yet, 2D IO components such as smartphones are restricted by their form factor, favoring device ergonomics and portability but limiting their potential output capabilities. By combining them with MR IO components (e.g., AR HWDs), we can “unchain” their capabilities, enabling entirely new interaction possibilities (e.g., offloading menu items [reipschlager2019designar, normand2018enlarging, brasier2021arenhanced]), thereby inching closer to a form of “universal interaction” [knierim2023universal] and ultimately, Weiser’s vision of ubiquitous computing [weiser1999computer]. While this can further expand the design space, we also recognize the need to establish guidelines to better understand the trade-offs and potential of each device (cf. [wang2023serious]).